You will receive $590,000 now, and you will not have to pay taxes on this amount. In addition, beginning one year from today, you will receive $260,000 each year for 30 years. The cash flows from this annuity will be taxed at 28 percent. Calculate PV assuming 7% discount rate
Answer #1

Current amount received = 590,000
After tax cash flow after 1 year for 30 years = 260,000*(1-28%) = 187200
PV of Cash Flows using annuity formula = 590,000 + 187200*(1-(1+7%)-30)/7% = 590,000 + 2,322,972.51= 2,912,972.51

Please Discuss in case of Doubt

Best of Luck. God Bless
Please Rate Well